OpenStudy (anonymous):

Write an equation for the tangent line to f(x) = (2x) - (3/x) at x=-2 I have around 9 more of these kinds of questions I need to solve. Please show some work. Thank you!

OpenStudy (zehanz):

You will have to take the derivative of f, and then calculate f'(-2). Do you know how to do this?

OpenStudy (anonymous):

Would the derivative be 2 - (3/x^2) ?

OpenStudy (zehanz):

Yes, it would! The equation you are looking for, has the form y = ax + b. Here, the slope a of the line is also f'(-2), so plug in this value.

OpenStudy (anonymous):

11/4 ?

OpenStudy (zehanz):

f'(-2) = 2 - 3/4 = 1 1/4, oh I see now what you mean. (1.25) So you have y = 1.25x + b. To find b, you will have to know a point that is on the line. That point is (-2, f(-2)), because that is the point where the tangent line "touches" the graph of f. Calculate f(-2). Then put the x and y in y=-2x + b to find b. Then you're done!

OpenStudy (zehanz):

I mean y = 1.25x + b

OpenStudy (anonymous):

f(-2) would be -2.5

OpenStudy (zehanz):

So (-2, -2.5) is on the line. y = 1.25x + b, so b = y - 1.25x. Set x = -2 and y = -2.5 to get b.

OpenStudy (anonymous):

I got b = -5

OpenStudy (anonymous):

answer choices are -10x-5y=-13, -11-4y = -12 11x-4y=-12 10x-5y=-13

OpenStudy (anonymous):

any idea which one?

OpenStudy (zehanz):

Oops. Error in f'(x). It must be 2 + 3/x^2 (see why?). So f'(-2)=2+3/4 = 2.75 y = 2.75x + b, b = y - 2.75x = -2.5 - 2.75*-2 = -2.5 + 5.5 = 3, so y = 2.75x + 3 or ( multiply by 4) 4y = 11x + 12, or 11x -4y = -12

OpenStudy (anonymous):

ah! thanks!